# L-MARS: Legal Multi-Agent Workflow with Orchestrated Reasoning and Agentic Search

### v2

@@ Abstract
We present L-MARS (Legal Multi-Agent Workflow with Orchestrated Reasoning and Agentic Search), a system that reduces hallucination and uncertainty in legal question answering through coordinated multi-agent reasoning and retrieval. Unlike single-pass RAG, L-MARS decomposes queries into subproblems, issues targeted searches across heterogeneous sources (Serper web, local RAG, CourtListener case law), and employs a Judge Agent to verify sufficiency, jurisdiction, and temporal validity before answer synthesis. This iterative reasoning–search–verification loop maintains coherence, filters noisy evidence, and grounds answers in authoritative law. We evaluated L-MARS on LegalSearchQA, a new benchmark of 200 up-to-date multiple choice legal questions in 2025. The results show that L-MARS substantially improves the accuracy of the factual data, reduces uncertainty, and achieves higher preference scores for the human and LLM judges. Our work demonstrates that multi-agent reasoning with agentic search is a scalable, reproducible blueprint for deploying LLMs in high-stakes domains requiring precise legal retrieval and deliberation. The code is available at: https://github.com/boqiny/L-MARS.
@@ 1 Introduction
Large language models (LLMs) such as GPT (gpt_4o_system_card, ), Claude (anthropic2024claude35sonnet, ), and Gemini (gemini2024arxiv, ) are increasingly applied to legal tasks such as case law retrieval (locke2022caselawretrievalproblems, ), statutory interpretation, and automated legal assistance(rincon2021legalclinic, ). Their ability to generate fluent context-sensitive responses makes them attractive for legal services, but direct application often results in hallucinations (ji2023survey, ; xu2025hallucinationinevitableinnatelimitation, ) or uncertainty (huang2024surveyuncertaintyestimationllms, )—confidently stated yet factually unsupported answers. In legal contexts, such errors carry significant real-world risk, as incorrect citations or outdated statutes can undermine credibility and thus negatively affect decision-making.
Two primary strategies are commonly used to mitigate hallucinations in domain-specific LLMs. Domain-specific fine-tuning(chalkidis2020legalbert, ) can improve performance in legal queries, but is costly and fragile. Legal systems evolve rapidly: new regulations are introduced, existing laws are amended, and obsolete provisions are repealed. Maintaining accuracy would require frequent retraining, which makes this approach impractical for continuously updated domains. Retrieval-Augmented Generation (RAG) (lewis2020rag, ) avoids the retraining burden by retrieving relevant documents from an external corpus at inference time, but its effectiveness depends heavily on retrieval accuracy. If key legal evidence is missed, the model may still hallucinate or produce incomplete reasoning.
To address these challenges, we introduce L-MARS, a system that integrates diverse retrieval sources with structured outputs and multi-turn model reasoning. L-MARS combines online search via the Serper API (serperapi, ) for up-to-date web information, offline retrieval using RAG over a curated local legal database, and case law retrieval through the CourtListener API (courtlistenerapi, ) for authoritative legal opinions. Unlike prior single-turn pipelines, L-MARS employs a multi-agent workflow that iteratively decomposes queries, executes targeted retrieval, and performs sufficiency checks before synthesizing final answers.
@@
L-MARS operates in two modes. In the Simple Mode, a single-turn pipeline retrieves relevant legal documents and uses a model with reasoning capacity to integrate them into the final answer. In the Multi-Turn Mode, a Query Agent refines the user query through clarifying sub-questions, a Judge Agent evaluates the sufficiency and relevance of retrieved evidence, and a Summary Agent synthesizes the final answer. This iterative loop enables the system to adaptively search, verify, and refine until an evidence sufficiency threshold is met.
By combining reasoning with agentic search, L-MARS achieves higher factual accuracy and significantly reduces uncertainty compared to the baseline of the GPT, Claude, and Gemini models. The framework also provides a flexible architecture that can switch between low-latency simple mode and high-accuracy multi-turn mode, offering a reproducible blueprint for high-stakes domains such as law.
Table 1: Examples of Model-Expressed Uncertainty in Legal QA Responses
Model Uncertainty Expressions in Pure LLM Responses for Legal Questions
Regulations may differ across various cities and provinces in China.
As of my last update in October 2023, public protest in China is heavily regulated…
Legal interpretations and implementations may vary between EU Member States.
This information is based on the legal framework as of 2023 and may be subject to change.
It is important to note that regulations are subject to change, especially with new laws or court rulings.
@@ 2 Related Work
Agentic RAG and Multi-Agent Framework
RAG grounds the model output in external evidence to reduce hallucination. Moving beyond single-turn pipelines, interleaving reasoning with retrieval (yao2023reactsynergizingreasoningacting, ; khattab2023, ) improves factual grounding and sample efficiency. Multi-agent frameworks (wu2023autogen, ; langgraph2024, ) operationalize role specialization (planner, retriever, verifier) and conversational coordination, enabling iterative retrieval, critique, and revision. Recent agentic RAG variants further combine planning and verification with retrieval to maintain long-horizon reasoning(singh2025agenticRAG, ).
Search Systems for LLMs
Long-horizon tool-using research requires agents that actively plan queries, browse, and synthesize sources. Search-o1 and WebThinker (searcho1, ; li2025webthinkerempoweringlargereasoning, ) integrates agentic search into large reasoning models to trigger retrieval at uncertainty points and reason over documents before injection into the chain. In production settings, OpenAI’s Deep Research (openai2025deepresearch, ) exposes an agentic, web-scale research loop designed for deeper, broader investigations. These systems complement “think-then-answer” reasoning models (openai2024o1arxiv, ; openai2025o3systemcard, ) by coupling deliberation with targeted information acquisition.
Reasoning Models and Techniques
Reasoning-oriented prompting has shown strong gains: chain-of-thought (wei2023chainofthoughtpromptingelicitsreasoning, ) prompting encourages models to externalize intermediate steps, while self-consistency reduces variance by aggregating diverse traces (wang2022selfconsistency, ). Structured interaction methods, including debate-style critique and verifier-based confidence estimation, further improve factuality and calibration (kadavath2022know, ). More recently, large reasoning models such as OpenAI’s o1(openai2024o1arxiv, ), Qwen-QwQ(qwq-32b-preview, ), and DeepSeek-R1(deepseek-r1, ) extend these ideas by scaling the deliberation at the test time, explicitly demonstrating multi-step reasoning traces across domains such as mathematics, code, and science.
@@
Legal AI has progressed through a series of domain-specific benchmarks and corpora. LegalBench(guha2023legalbench, ) measures practical legal reasoning for diverse tasks ; LexGLUE(chalkidis2022lexglue, ) standardizes the evaluation across multi-task legal NLU; and Pile-of-Law(henderson2022pileoflaw, ) provides a large and responsibly curated pretraining corpus. More recent work has introduced reasoning and retrieval focused evaluations: Reasoning-Focused Legal Retrieval Benchmark(Zheng_2025, ) introduces Bar Exam QA and Housing Statute QA to capture realistic legal RAG settings, while LEXam (fan2025lexambenchmarkinglegalreasoning, ) benchmarks long-form, multi-step legal reasoning with law exam questions in English and German. Our work builds on this trajectory by combining retrieval with agentic multi-turn reasoning and domain-specific verification, complementing these benchmarks with a system architecture explicitly designed to reduce hallucination and uncertainty.
@@ 3.1 Problem Formation
We formalize legal question answering as generating both a reasoning chain $R$ and a final answer $a$ given a user query $q$, external retrieval results $D$, and optional user clarifications $U$. Each query state is represented as $(q,U,D)$, which evolves iteratively through the workflow. At each step, the system decides whether (i) expand $D$ by issuing new search queries, or (ii) terminate with a sufficient answer. Formally, the objective is to learn a mapping
@@ 3.2 Workflow and Agents Overview
L-MARS is a multi-agent workflow implemented in LangGraph with structured output validation and tool use. The system employs a directed acyclic graph (DAG) architecture in which agents are represented as nodes and the control flow is represented as edges with conditional routing. The workflow maintains a centralized WorkflowState that tracks: (i) the evolving query representation, (ii) the accumulated search results with metadata, (iii) the iteration history and refinement notes, and (iv) the intermediate agent outputs. State transitions are deterministic and type-checked, ensuring reproducibility across runs.
@@
The Query agent parses the user’s question into a structured query result (issue type, key entities, time window, jurisdiction, and initial search intents). The Search Agent executes tool calls to the retrieval backends and returns normalized SearchResult objects (title, url, snippet, optional full content, source type). The Judge Agent performs evidence sufficiency checks and an explicit checklist (factual support, jurisdiction match, temporal specificity, and contradiction analysis). Summary Agent composes the final response with citations and rationale. Detailed agent implementations can be found in the Appendix A.
@@ 3.3 Agentic Search
We adopt an agentic retrieval policy inspired by Search-o1(searcho1, ), interleaving targeted web queries with stepwise reasoning rather than performing a single problem-oriented fetch. Concretely, the Search Agent triggers retrieval when either (i) the Judge Agent flags a knowledge gap (e.g. missing statute/case authority, date or jurisdiction ambiguity), or (ii) the query classifier predicts that external evidence will reduce answer uncertainty. Once activated, the agent executes one of two complementary search modes against Google Serper, as shown in Table 2 :
@@
Enhanced Search (content extraction) uses the top-$m$ URLs selected by the agent (default m=3m{=}3), then fetches page contents and PDFs with robust headers and timeouts. HTML is cleaned with BeautifulSoup; PDFs are parsed with pdfplumber. To avoid flooding the context, we perform a snippet-anchored extraction: given the search snippet, we locate the best-matching sentence in the full text using a token-level $F_{1}$ overlap and return a fixed window of the surrounding context (2.5k chars). This preserves local evidence while discarding boilerplate and navigation chrome. The extractor also normalizes whitespace, strips punctuation, and truncates long pages to a hard cap to guaranty bounded latency.
@@
We also maintain a local retrieval index over user-provided documents. Markdown files are segmented into overlapping windows of 500 characters with 100-character stride and indexed using the BM25 ranking function (robertson2009bm25, ). At query time, the retriever returns the top-$k$ segments (default $k=5$), enriched with document titles and section headers to preserve context. The indexing process is dynamic: new files added to the inputs directory are automatically detected and incorporated without requiring a system restart.
@@ 3.4 L-MARS Operating Mode
Algorithm 1 L-MARS Multi-Turn Workflow
1:Input: user query $q$, max iterations $M$
[…]
4:Receive user responses $U$;  $Q\leftarrow\textsc{Incorporate}(Q,U)$
[…]
6:for $i=1$ to $M$ do
7:  $S\leftarrow\textsc{GenSearchQueries}(Q,R,E)$
8:  $R_{i}\leftarrow\textsc{DeepSearch}(S,k{=}3,\text{with\_content}{=}\text{True})$
9:  $(\text{suff},\text{notes})\leftarrow\textsc{JudgeAgent}(R\cup R_{i},T{=}0)$ $\triangleright$ CoT, quality, date/jurisdiction, contradiction
10:  $R\leftarrow R\cup R_{i}$;  $E\leftarrow E\cup\text{notes}$
11:  if $\text{suff}=\textsc{Sufficient}$ then
12:   break
13:  else
14:   $Q\leftarrow\textsc{Refine}(Q,\text{notes},U)$
15:  end if
16:end for
[…]
18:return $(a)$
Both modes share the same retrieval backends and dual evaluation stack. The key distinction lies in the graph topology: Simple Mode follows a linear path START $\to$ Query $\to$ Search $\to$ Answer $\to$ END, while Multi-Turn Mode includes a conditional loop with the Judge Agent determining whether to continue retrieval or proceed to answer generation based on evidence sufficiency criteria.
@@
By default, the Judge Agent is instantiated with GPT-o3 for stronger reasoning ability, and runs with temperature =0=0 to reduce variance and improve reproducibility. Its prompt elicits three outputs: (i) a chain-of-thought style rationale, (ii) a structured checklist decision on sufficiency, and (iii) missing-evidence directives that guide the next refinement queries. The loop ends once the decision is made Sufficient or when the maximum iteration limit is reached. In practice, we found the Judge Agent especially useful when search results included unofficial forums (e.g., Reddit, Quora) instead of government domains; in such cases it flagged insufficiency and steered subsequent searches toward more compelling, authoritative sources such as government websites and official regulations.
@@ 4.1 Tasks and Datasets
We introduce LegalSearchQA, a benchmark of 200 legal questions designed to target post-training knowledge gaps. Unlike prior benchmarks such as LegalBench or CUAD(hendrycks2021cuadexpertannotatednlpdataset, ), which evaluate reasoning over pre-provided context, UncertainLegalQA tests the end-to-end ability of a system to retrieve and reason over external legal sources—an essential capability for real-world deployment. To avoid reliance on static pre-training knowledge, all questions explicitly reference the legal status as of 2025, requiring retrieval of up-to-date authorities.
The dataset spans multiple domains, with a strong emphasis on federal executive orders (60%), along with coverage of tax provisions (10%), immigration (10%), federal tax inflation updates (10%), housing reforms (5%) and technology law (5%). This distribution reflects a focus on recent federal regulatory actions while still incorporating broader legislative and policy areas. Jurisdictional coverage includes both US federal law and major state-level developments, emphasizing contexts of regulatory flux, pending rulemaking, and areas where conflicting authorities require active retrieval. The questions are further categorized into requirement checks (48%), permissibility assessments (12%), definitions and scope (10%), policy intent (8%), timeline queries (8%), institutional authority (6%), legal authority references (4%) and eligibility thresholds (4%).
Figure 2: Composition of the LegalSearchQA dataset.
4.2 Evaluation Metrics
We employ a comprehensive evaluation framework that combines automated metrics with LLM-based judgment to capture both quantitative and qualitative aspects of LegalSearchQA.
Accuracy
measures the fraction of multiple choice questions answered correctly against expert-annotated ground truth, reflecting the model’s ability to select the legally correct option.
U-Score
We propose U-Score, a rule-based metric for the evaluation of legal QA systems based on uncertainty. It ranges from 0 to 1 (lower is better) and captures reliability across five complementary dimensions: (i) hedging cues: frequency of uncertainty markers such as “may” or “could”; (ii) temporal vagueness: imprecise or underspecified references to time (e.g., “recently” instead of a year); (iii) citation sufficiency and authority: rewarding the use of primary legal sources (statutes, regulations, case law) over secondary commentary; (iv) jurisdictional specificity: explicit recognition of the relevant legal jurisdiction (e.g., U.S. federal vs. California state law); and (v) decisiveness: whether the answer delivers a clear conclusion rather than remaining equivocal. These components are linearly aggregated with weights:
$\textbf{U-Score}=0.25\,H+0.20\,T+0.25\,(1-C)+0.15\,(1-J)+0.15\,(1-D),$ (1)
where $H,T,C,J,D\in[0,1]$ denote hedging, temporal vagueness, citation sufficiency, jurisdictional specificity, and decisiveness, respectively.
LLM-as-Judge
complements these metrics with qualitative ratings using GPT-o3, evaluating answers along four dimensions: factual accuracy (consistency with current law), evidence grounding (relevance and credibility of cited sources), clarity of reasoning (logical coherence and completeness) and uncertainty calibration (appropriate hedging in genuinely ambiguous areas). Each response is assigned an overall rating of low, moderate, or high, reflecting the holistic quality of legal reasoning. Together, this dual evaluation framework captures both surface-level correctness and deeper reasoning quality.

### v3

@@ Abstract
We present L-MARS (Legal Multi-Agent Workflow with Orchestrated Reasoning and Agentic Search), a multi-agent retrieval framework for grounded legal question answering that decomposes queries into structured sub-problems, retrieves evidence via agentic web search, filters results through a verification agent, and synthesizes cited answers. Existing legal QA benchmarks test either closed-book reasoning or retrieval over fixed corpora, but neither captures scenarios requiring current legal information. We introduce LegalSearchQA, a 50-question benchmark across five legal domains whose answers depend on recent developments that post-date model training data. L-MARS achieves 96.0% accuracy in LegalSearchQA, a 38.0% improvement over zero-shot (58.0%), while chain-of-thought prompting degrades performance to 30.0%. On Bar Exam QA [28], a reasoning-focused benchmark of 594 bar examination questions, retrieval provides negligible gains (+0.7 percentage points), consistent with prior findings. These results show that agentic retrieval dramatically improves legal QA when tasks require up-to-date factual knowledge, but the benefit is benchmark-dependent, underscoring the need for retrieval-focused evaluation. Code and data: https://github.com/boqiny/L-MARS.
@@ 1 Introduction
Large language models (LLMs) such as GPT [1], Claude [3], and Gemini [4] are increasingly applied to legal tasks including case law retrieval [26], statutory interpretation, and automated legal assistance [2]. However, direct application often produces hallucinations [5, 27] where the model states confidently incorrect legal facts. In legal contexts such errors carry real-world risk, as wrong citations or outdated statutes can undermine credibility and affect decision-making.
Domain-specific fine-tuning [6] can help but requires frequent retraining as laws change, making it impractical for continuously evolving domains. Retrieval-Augmented Generation (RAG) [7] avoids retraining by fetching relevant documents at inference time, but its effectiveness depends on retrieval quality. Recent work has shown that on reasoning-intensive legal benchmarks, standard retrieval pipelines barely improve downstream QA accuracy. On Bar Exam QA, even the best traditional retriever improves GPT-4o-mini accuracy by only 0.5 percentage points over zero-shot [28].
We introduce L-MARS, a multi-agent retrieval framework that addresses these challenges through structured agent orchestration. L-MARS combines online search via the Serper API [8], offline retrieval using BM25 over local documents, and case law retrieval through the CourtListener API [9]. Unlike naive retrieve-and-generate pipelines, L-MARS employs a multi-agent workflow that decomposes queries into structured search intents, executes targeted retrieval, filters evidence through a Judge Agent, and synthesizes grounded answers.
@@
L-MARS operates in two modes. In Simple Mode, a single-pass pipeline decomposes the query, retrieves and filters evidence, and synthesizes a grounded answer. In Multi-Turn Mode, an iterative loop enables the system to adaptively search, verify, and refine until an evidence sufficiency threshold is met. In this work we evaluate Simple Mode to isolate the contribution of agent orchestration from iterative refinement.
In legal practice, outdated information can lead to materially incorrect advice. Our results suggest that retrieval is essential for legal QA tasks that depend on current facts, and may be important for reliable deployment in domains where knowledge changes rapidly.
Our contributions are threefold. First, we introduce LegalSearchQA, a benchmark for legal QA requiring post-training, time-sensitive knowledge. Second, we present L-MARS, a multi-agent retrieval framework for grounded legal question answering. Third, we show empirically that retrieval yields large gains only in retrieval-dependent settings, while chain-of-thought reasoning can degrade performance when parametric knowledge is outdated. All prompts, evaluation scripts, and raw outputs are released to ensure full reproducibility.
@@ 2 Related Work
Agentic RAG and Search Systems
RAG grounds model output in external evidence to reduce hallucination [7], and interleaving reasoning with retrieval [30, 10] further improves factual grounding. Multi-agent frameworks [11, 12] enable role specialization and iterative critique, with recent agentic RAG variants combining planning, verification, and retrieval for long-horizon tasks [31, 33]. On the search side, Search-o1 and WebThinker [13, 14] integrate agentic search into reasoning models, while OpenAI’s Deep Research [15] exposes a production-grade web-scale research loop. These systems complement reasoning models [16, 17] that scale deliberation at test time through chain-of-thought prompting [18], self-consistency [19], verifier-based calibration [32], and dedicated reasoning architectures [20, 21].
@@
Legal AI evaluation spans benchmarks for reasoning [24], multi-task NLU [22], Chinese legal knowledge [34], and pretraining data [23]. More recent work targets retrieval and reasoning specifically: Bar Exam QA [28] captures realistic legal RAG settings where standard retrieval fails, and LEXam [29] benchmarks multi-step legal reasoning across languages. On the systems side, ChatLaw [35] explores multi-agent approaches for legal tasks. We note that many LegalBench [24] subtasks are now saturated, with GPT-4o-mini achieving 100% on the rule_qa, limiting their discriminative value for knowledge-intensive legal QA.
@@ 3.1 Problem Formation
We formalize legal question answering as generating both a reasoning chain $R$ and a final answer $a$ given a user query $q$, external retrieval results $D$, and optional user clarifications $U$. Each query state is represented as $(q,U,D)$, which evolves iteratively through the workflow. At each step, the system decides whether (i) expand $D$ by issuing new search queries, or (ii) terminate with a sufficient answer. Formally, the system computes a mapping
@@ 3.2 Workflow and Agents Overview
L-MARS is a multi-agent workflow implemented in LangGraph with structured output validation and tool use. The system employs a directed graph in which agents are represented as nodes and the control flow is represented as edges with conditional routing. The workflow maintains a centralized WorkflowState that tracks: (i) the evolving query representation, (ii) the accumulated search results with metadata, (iii) the iteration history and refinement notes, and (iv) the intermediate agent outputs. State transitions are deterministic and type-checked, ensuring reproducibility across runs.
@@
The Query agent parses the user’s question into a structured query result. The Search Agent executes tool calls to the retrieval backends and returns normalized SearchResult objects (title, url, snippet, optional full content, source type). The Judge Agent performs evidence sufficiency checks and an explicit checklist (factual support, jurisdiction match, temporal specificity, and contradiction analysis). Summary Agent composes the final response with citations and rationale. Detailed agent implementations can be found in the Appendix A.
@@ 3.3 Agentic Search
We adopt an agentic retrieval policy inspired by Search-o1[13], interleaving targeted web queries with stepwise reasoning rather than performing a single problem-oriented fetch. Concretely, the Search Agent triggers retrieval when either (i) the Judge Agent flags a knowledge gap (e.g. missing statute/case authority, date or jurisdiction ambiguity), or (ii) the query classifier predicts that external evidence will reduce answer uncertainty. Once activated, the agent executes one of two complementary search modes against Google Serper, as shown in Table 1 :
@@
Enhanced Search (content extraction) uses the top-$m$ URLs selected by the agent (default m=3m{=}3), then fetches page contents and PDFs with robust headers and timeouts. HTML is cleaned with BeautifulSoup; PDFs are parsed with pdfplumber. To avoid flooding the context, we perform a snippet-anchored extraction: given the search snippet, we locate the best-matching sentence in the full text using a token-level $F_{1}$ overlap and return a fixed window of the surrounding context (2.5k chars). This preserves local evidence while discarding boilerplate and navigation chrome. The extractor also normalizes whitespace, strips punctuation, and truncates long pages to a hard cap to guarantee bounded latency.
@@
We also maintain a local retrieval index over user-provided documents. Markdown files are segmented into overlapping windows of 500 characters with 100-character stride and indexed using the BM25 ranking function [25]. At query time, the retriever returns the top-$k$ segments (default $k=5$), enriched with document titles and section headers to preserve context. The indexing process is dynamic: new files added to the inputs directory are automatically detected and incorporated without requiring a system restart.
@@ 3.4 L-MARS Operating Mode
The Judge Agent runs with temperature =0=0 to reduce variance and improve reproducibility. Its prompt elicits three outputs: (i) a chain-of-thought style rationale, (ii) a structured checklist decision on sufficiency, and (iii) missing-evidence directives that guide the next refinement queries. The loop ends once the decision is made Sufficient or when the maximum iteration limit is reached. In practice, we found the Judge Agent especially useful when search results included unofficial forums (e.g., Reddit, Quora) instead of government domains; in such cases it flagged insufficiency and steered subsequent searches toward more compelling, authoritative sources such as government websites and official regulations.
We describe both operating modes and all three retrieval backends for completeness, as they form the full L-MARS system available in our codebase. In our evaluation we use only Simple Mode with Serper web search, since this minimal configuration already achieves 96% accuracy on LegalSearchQA. Local RAG indexing and CourtListener case law retrieval are available as optional backends for practitioners working with proprietary document collections or case-specific research, but are not exercised in our benchmark evaluation.
@@ 4.1 Tasks and Datasets
LegalSearchQA
Existing legal QA benchmarks test either closed-book reasoning on static exam questions or retrieval over fixed corpora. Neither reflects the real-world setting where a legal professional needs current information, such as recently enacted statutes, updated regulatory thresholds, or new court decisions, that was not available when the model was trained. To address this gap we introduce LegalSearchQA, a 50-question multiple-choice benchmark designed to evaluate retrieval-augmented legal QA systems on current legal knowledge.
Each question has four answer choices and exactly one correct answer. Questions span five legal domains: Tax, Corporate & Financial Regulation (13), Labor & Employment (11), Immigration (9), Technology & Privacy (9), and Criminal, Drug & State Law (8). The benchmark targets three categories of legal knowledge: status questions about the current state of laws, threshold questions about specific numerical values such as dollar amounts or deadlines, and requirement questions about what rules mandate. All questions were authored by the paper’s authors and verified against primary legal authorities such as IRS.gov, USCIS.gov, SEC.gov, and Supreme Court opinions during March 2026. By construction, correct answers require information from 2024 to 2026 that is unlikely to appear in model training data, making web retrieval essential. To reduce author bias, each question is grounded in a verifiable primary legal source and the task is purely factual retrieval rather than subjective interpretation, limiting degrees of freedom in answer labeling. While the absolute accuracy of 96% is high, this reflects the benchmark’s design: questions are intentionally selected to isolate retrieval-dependent knowledge where access to current information provides a clear advantage.
To further reduce potential bias, we release all questions with source citations, timestamps, and verification notes. Each question is constructed to have a single unambiguous factual answer grounded in primary legal sources (e.g., IRS, SEC, Federal Register). Because the task is factual rather than interpretive, correctness can be independently verified without subjective judgment. The complete dataset is available at https://huggingface.co/datasets/boqiny/LegalSearchQA.
Bar Exam QA
We additionally evaluate on Bar Exam QA [28], a legal QA benchmark drawn from the Reasoning-Focused Legal Retrieval Benchmark. The dataset consists of multistate bar examination questions that test legal reasoning across core areas of U.S. law. We use the publicly released Historical MBE subset of 594 questions with subject labels spanning six domains: Contracts (113), Torts (112), Constitutional Law (95), Evidence (93), Real Property (92), and Criminal Law (89).
The authors of Bar Exam QA showed that standard retrieval pipelines achieve very low recall on this benchmark and their best retriever improves GPT-4o-mini accuracy by only 0.5 percentage points over zero-shot. We include this benchmark to confirm that retrieval provides negligible gains on reasoning-focused tasks. We note that the Historical MBE subset overlaps with the MMLU auxiliary training set [36] and may inflate zero-shot baselines through memorization; we mitigate this by focusing primary claims on LegalSearchQA, where memorization is not a factor.
4.2 Baselines and Experimental Conditions
We compare three conditions, all using GPT-4o-mini as the backbone model:
Zero-Shot
The model answers directly with no retrieval and no explicit reasoning instructions. The prompt presents the question and four choices, instructing the model to select a single answer letter.
Chain-of-Thought (CoT)
The model is prompted to reason step-by-step: identify the legal issue, recall the relevant rule, apply it to the facts, and eliminate incorrect options before selecting an answer. No external retrieval is provided.
L-MARS (Simple Mode)
The full L-MARS pipeline in single-pass mode: the Query Agent decomposes the question into structured search intents, the Search Agent retrieves from Serper with enhanced content extraction, and the Summary Agent synthesizes the final answer with citations. All agents use GPT-4o-mini with T=0T{=}0.
This experimental ladder isolates specific contributions: Zero-Shot to CoT tests whether reasoning alone helps, and CoT to L-MARS tests whether agentic retrieval adds value beyond reasoning. We leave evaluation of the Multi-Turn Mode to future work, as it targets open-ended legal research questions where a single retrieval pass may not surface sufficient evidence, a setting not captured by our current multiple-choice benchmarks.
While L-MARS includes multiple components, our evaluation isolates the contribution of retrieval itself by comparing zero-shot against the full pipeline. Since Simple Mode excludes iterative refinement and Judge Agent loops, the observed improvements can be attributed primarily to structured web retrieval rather than multi-turn reasoning.